OpenStudy (anonymous):

Find all the values of x, between 0 and 360 which satisfy the equation 3 sin x= 2 + 1/sin x

OpenStudy (gir_lover_♥):

6cos^2x+5sinx=2 6(1-sin^2x)+5sinx-2=0 6-6sin^2x+5sinx-2=0 -6sin^2x+5sinx+4=0 6sin^2x-5sinx-4=0 solve by factoring this quadratic equation (3sinx-4)(2sinx+1)=0 3sinx-4=0 3sinx=4 sinx=4/3 (reject,sinx<=1)

OpenStudy (gir_lover_♥):

The sin and x are suppose to be separated sawry bout that

OpenStudy (anonymous):

it's fine.. thank u

OpenStudy (gir_lover_♥):

Ok good